# Multiscale Representation Enhanced Temporal Flow Fusion Model for Long-Term Workload Forecasting

### v1

@@ Abstract.
time series, workload forecasting
@@ 1. Introduction
  • Challenge 1: The multi-periodic, non-stationary, and long-term dependencies. In long-term forecasting, workloads time series data is often high-frequency and non-stationary with complex multi-periodic characteristics, such as hourly, daily, and weekly cycles. Furthermore, these data exhibit long-term dependent properties. Hence, capturing the multi-periodic and long-term characteristics of time series data is critical in addressing the challenges of long-term forecasting.
@@ 4.1. Benchmarks
For long-term forecasting, we conduct the experiments on 9 well-established benchmarks: ETT datasets (including 4 subsets: ETTh1, ETTh2, ETTm1, ETTm2) (Zhou et al., 2021), Solar-Energy(Lai et al., 2018), Weather, Electricity, and Traffic (Wu et al., 2021), as well as the large Service-Workload datasets. More information is provided in the Appendix.
@@ 4.2. Baselines
We compared our method with $10$ advanced baselines. These methods can be divided into two categories: end-to-end forecasting models (including PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Non-stationary Transformer (Liu, 2022), Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021)) and time series representation models (including TimeMAE (Cheng et al., 2023), LaST (Wang et al., 2022a), TST(Zerveas et al., 2021), COST (Woo et al., 2022), TS2VEC (Yue et al., 2022)). The detailed descriptions and implementations of experiments are provided in the Appendix. Regarding metrics, we utilize the mean square error (MSE) and mean absolute error (MAE) for long-term forecasting.
@@ 4.4. Model Analysis
Figure 4. Visualization of TS representation of two weeks on workload dataset. Changes in daily and weekly periods are clearly indicated.
@@
Figure 8. The performance of predictive autoscaling deployed in the production environment. Blue vertical dotted line indicates the launch of predictive autoscaling based on workload forecasting. Green frame highlights the comparison of the metrics under the same workload before and after enabling predictive autoscaling.
@@ 5. Application
Alipay, a leading global mobile payment company, has established its presence in various domains including digital life and digital finance (Chu et al., 2022). With an expansive cloud computing infrastructure supporting countless microservices, efficient allocation of cloud resources is critical for Alipay’s cluster management. Alipay relies on its proprietary predictive autoscaling technology as the fundamental component of its cloud resource management system. The accurate prediction of future workloads for each microservice, specifically in terms of requests per second (RPS), is crucial for the effective scaling of server resources. Presently, our proposed approach has been extensively deployed in Alipay’s production cloud environment, achieving remarkable results. The outcomes, as demonstrated in Table 1, indicate a significant improvement of 26% in comparison to the state-of-the-art (SOTA) methods in the real-world environment. Figure 8 demonstrates the utilization of our autoscaling method based on workload forecasting, which enables the Alipay cloud resource management system to anticipate changes in future workloads through long-term forecasting when traffic fluctuates. This facilitates the scaling of computing resources to allocate those appropriate for the current workload. Upon activating our prediction technique, the number of pods (containers that host microservices) utilized by the microservices has significantly decreased from 1500 to less than 500 (Reduced resource consumption by 67%), leading to a substantial improvement in resource utilization efficiency.

### v2

@@ Abstract.
time series, workload forecasting, multiscale representation
@@ 1. Introduction
  • Challenge 1: The multi-periodic, non-stationary, and long-term dependencies. In long-term forecasting, workloads time series data is usually high-frequency and non-stationary with complex multi-periodic characteristics, such as hourly, daily, and weekly cycles. Furthermore, these data exhibit long-term dependent properties. Hence, capturing the multi-periodic and long-term characteristics of time series data is critical in addressing the challenges of long-term forecasting.
@@ 4.1. Benchmarks
For long-term forecasting, we conduct the experiments on 9 well-established benchmarks: ETT datasets (including 4 subsets: ETTh1, ETTh2, ETTm1, ETTm2) (Zhou et al., 2021), Solar-Energy(Lai et al., 2018), Weather, Electricity, and Traffic (Wu et al., 2021), as well as the large Service-Workload datasets.
@@ 4.2. Baselines
We compared our method with $10$ advanced baselines. These methods can be divided into two categories: end-to-end forecasting models (including PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Non-stationary Transformer (Liu, 2022), Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021)) and time series representation models (including TimeMAE (Cheng et al., 2023), LaST (Wang et al., 2022a), TST(Zerveas et al., 2021), COST (Woo et al., 2022), TS2VEC (Yue et al., 2022)). Regarding metrics, we utilize the mean square error (MSE) and mean absolute error (MAE) for long-term forecasting.
@@ 5. Application
Alipay, a leading global mobile payment company, has established its presence in various domains including digital life and digital finance (Chu et al., 2022). With an expansive cloud computing infrastructure supporting countless microservices, efficient allocation of cloud resources is critical for Alipay’s cluster management. Alipay relies on its proprietary predictive autoscaling technology as the fundamental component of its cloud resource management system. The accurate prediction of future workloads for each microservice, specifically in terms of requests per second (RPS), is crucial for the effective scaling of server resources. Presently, our proposed approach has been extensively deployed in Alipay’s production cloud environment, achieving remarkable results. The outcomes, as demonstrated in Table 1, indicate a significant improvement of 26% in comparison to the state-of-the-art (SOTA) methods in the real-world environment. It demonstrates the utilization of our autoscaling method based on workload forecasting, which enables the Alipay cloud resource management system to anticipate changes in future workloads through long-term forecasting when traffic fluctuates. This facilitates the scaling of computing resources to allocate those appropriate for the current workload. Upon activating our prediction technique, the number of pods (containers that host microservices) utilized by the microservices has significantly decreased from 1500 to less than 500 (Reduced resource consumption by 67%), leading to a substantial improvement in resource utilization efficiency.